Main INDEX
Monthly INDEX
PREV
NEXT
User name R. Michaels
Log entry time 17:21:04 on November 1,2000
Entry number 51352
keyword=trigger checkout
After rolling the L-arm spectrometer back in place, we
reconnected the cables and adjusted timing. Note: T0
changed, need recalibration. The coincidence window width
should be about 105 nsec (95 nsec width on R-arm, 20 nsec
width after adjustment from L-arm minus a minimum overlap
of 5 nsec on each side). It would be wise to throw out all but
the middle 80 nsec of this range, due to retiming window effects.
Only one trigger setting is possible, as evident when you run ``trigsetup''.
T2 (R-arm) uses Cerenkov, T4 (L-arm) is ``or'' of S1,S2.
Regarding new trigger 6 (something from the calorimeter): We
did NOT plug it into scalers. This could be done with a cost of 14 nsec
delay of trigger due to fanout + cabling, but it was decided this extra
delay was not worth it. I guess its a test trigger that will normally be off.
Remember that any trigger not scaled can harm normalization (don't care ?).
Bottom line, we're ready to run.